# Particle-Grid Neural Dynamics for Learning Deformable Object Models from RGB-D Videos

### v1

@@ Abstract
Modeling the dynamics of deformable objects is challenging due to their diverse physical properties and the difficulty of estimating states from limited visual information. We address these challenges with a neural dynamics framework that combines object particles and spatial grids in a hybrid representation. Our particle-grid model captures global shape and motion information while predicting dense particle movements, enabling the modeling of objects with varied shapes and materials. Particles represent object shapes, while the spatial grid discretizes the 3D space to ensure spatial continuity and enhance learning efficiency. Coupled with Gaussian Splattings for visual rendering, our framework achieves a fully learning-based digital twin of deformable objects and generates 3D action-conditioned videos. Through experiments, we demonstrate that our model learns the dynamics of diverse objects—such as ropes, cloths, stuffed animals, and paper bags—from sparse-view RGB-D recordings of robot-object interactions, while also generalizing at the category level to unseen instances. Our approach outperforms state-of-the-art learning-based and physics-based simulators, particularly in scenarios with limited camera views. Furthermore, we showcase the utility of our learned models in model-based planning, enabling goal-conditioned object manipulation across a range of tasks. The project page is available at https://kywind.github.io/pgnd.
{strip}[Uncaptioned image]
Figure 1: Modeling deformable objects from RGB-D videos presents a significant challenge due to occlusions and complex physical interactions. Our Particle-Grid Neural Dynamics framework learns the behavior of deformable objects directly from real-world observations. To train the model, we introduce a novel dense 3D tracking method that leverages foundational vision models for video tracking. The trained model predicts the motion of dense particles under robot-object interactions. We demonstrate the ability of Particle-Grid Neural Dynamics to model complex interactions across a diverse set of objects, including ropes, cloth, plush toy, box, and bread.
@@ I Introduction
Learning predictive models is crucial for a wide range of robotic tasks. In deformable object manipulation, an accurate predictive object dynamics model enables model-based planning, policy evaluation, and real-to-sim asset generation. However, developing dynamics models for deformable objects that are both accurate and generalizable remains a significant challenge. For example, physics-based simulators [12, 31] often struggle to generalize to the real world due to the inherent sim-to-real gap and the difficulties of system identification and state estimation. Meanwhile, video-based predictive models [9, 56] are computationally expensive, lack 3D spatial understanding, and are highly sensitive to viewpoint and appearance changes.
@@
Recent advancements in deformable object modeling have focused on learning dynamics models for particles that encode 3D geometric information directly obtained from RGB-D cameras. The current state-of-the-art methods in this area utilize Graph Neural Networks (GNNs), where particle sets are modeled as spatial adjacency graphs, and message passing is performed to aggregate information and generate motion predictions [48, 58]. However, these methods face significant challenges: the effectiveness of message passing is highly sensitive to the spatial distribution and connectivity of the graph nodes, making them vulnerable to partial observations. Additionally, insufficient message-passing steps may fail to capture global information, while excessive steps can lead to overly smoothed predictions. As a result, these approaches are often limited to relatively simple simulated environments or real-world objects with easily perceivable geometries, where graph construction is straightforward with nearest neighbors.
@@
To address these limitations, we introduce a novel class of dynamic models called particle-grid neural dynamics. This model leverages a hybrid representation that combines object particles with fixed spatial grids. It takes the kinematic states of the particles as input and predicts a spatial velocity field at fixed grid points. A global point cloud encoder is employed to capture comprehensive information from the particle set, effectively overcoming the connectivity challenges commonly faced by GNNs and enhancing the model’s robustness to incomplete observations. Additionally, this encoder allows for the processing of denser particle sets compared to sparse graph nodes, thereby enriching the model with finer geometric details. The grid representation further acts as a regularizer, ensuring spatial continuity in velocity predictions while also reducing computational costs when handling large numbers of particles. By combining object particles with spatial grids, our framework parameterizes dynamics in both Lagrangian and Eulerian coordinates, drawing an analogy to physics-based deformable object simulation methods [44, 12]. By utilizing neural networks as message integrators to bridge particle and grid representations, our model handles a diverse range of materials without requiring material-specific physical laws, and can operate effectively with only partial observations as input.
@@
Notably, our model is trained purely on RGB-D videos of robots interacting with objects through diverse and non-task-specific behaviors. To extract object particles from raw recordings, we have developed a 3D fusion and tracking framework. This framework utilizes foundational vision models [18, 39, 15] to estimate segmentation masks and pixel tracks, which are then fused in 3D to generate persistent, dense 3D tracks that serve as training data for the neural dynamics model.
@@
In our experiments, we demonstrate that the proposed particle-grid neural dynamics model can effectively simulate a diverse range of challenging deformable objects, including ropes, cloth, plush toys, boxes, paper bags, and bread. We compare our model with existing state-of-the-art approaches and show that it consistently outperforms the baselines in prediction accuracy. Additionally, we highlight that our dynamics model seamlessly integrates with 3D appearance reconstruction methods such as 3D Gaussian Splatting (3DGS) [16], enabling the fusion of both techniques to generate highly realistic renderings of predictions. This improved accuracy in dynamics prediction directly translates to higher-fidelity renderings compared to existing methods. To further validate its robustness, we conduct experiments that demonstrate the model’s effectiveness under sparse-view conditions. Finally, we showcase the model’s applicability to deformable object manipulation by incorporating it into a Model Predictive Control (MPC) framework, underscoring its potential for real-world robotic applications.
@@ II Related Work
The simulation of deformable objects is essential for advancing both the modeling and robotic manipulation of soft, flexible materials. Researchers have introduced various analytical physics-based approaches, including, but not limited to, the mass-spring system [24, 20], the Finite Element Method (FEM) [10, 33, 7], the Discrete Elastic Rods (DER) [5, 6], the Position-Based Dynamics (PBD) [31, 21, 23], and the Material Point Method (MPM) [44, 12, 13, 29]. However, real-world analytical deformable modeling remains challenging due to the difficulty in property identification and state estimation. While our method uses a hybrid particle-grid representation similar to MPM, we leverage neural networks as message integrators and reduce dependence on full-state information as input. This enhances robustness to partial observations and enables the handling of a wide variety of deformable objects without requiring predefined material-specific constitutive laws. Recently, the emergence of 3D Gaussian Splatting (3DGS)  [16, 28, 8] has enabled the fusion of Gaussian Splatting reconstructions with physics-based models to simulate the dynamics of deformable objects [53, 37, 60, 59, 1]. In our particle-grid neural dynamics model, the particle motion predictions can also be integrated with Gaussian Splatting reconstructions, facilitating the simultaneous modeling and rendering of deformable objects, purely learned from real data.
@@
Learning-based dynamics models, which use deep neural networks to model the future evolution of dynamical systems, have demonstrated effectiveness across various robotic tasks [11, 55, 17, 61, 54, 4]. Among these learning-based methods, Graph-Based Neural Dynamics (GBND) have shown great promise, as they explicitly model spatial relational biases within complex physical systems [19, 34, 40, 3, 48]. Previous research has investigated the use of GBND across a range of material types, such as rigid bodies [14, 25, 2], plasticine [42, 43], fabrics [22, 26, 35, 27], ropes [46], and granular piles [47, 41]. Beyond simulation and single-material scenarios, GBND has also demonstrated flexibility and generalization in modeling diverse materials using a unified framework [57, 58]. However, these approaches often operate on spatially sparse graph vertices, rely on expert knowledge to determine graph connectivity, and do not consider partial observations. For learning dense particle dynamics, Whitney et al. [49] propose transformer-based backbones for higher computational efficiency. However, their work mainly focuses on modeling rigid objects for grasping and pushing tasks. In contrast, our work emphasizes deformable object modeling using a hybrid particle-grid neural dynamics framework, achieving dense particle prediction while remaining robust to incomplete observations and flexible for diverse types of deformable objects with distinct physical properties.
@@ III-A4 The Particle-Grid Model
Learning the neural dynamics parameters $\theta$ with an end-to-end neural network usually leads to unstable predictions due to accumulative error. Motivated by the use of hybrid Lagrangian-Eulerian representations in MPM [44], our model also utilizes a hybrid particle-grid representation to inject inductive bias related to spatial continuity and local information integration. Specifically, we define a uniformly distributed grid in the space as
@@
[…]
@@ III-B1 Point Encoder
We use PointNet [36] as the encoder for its efficiency and strong performance in extracting 3D point features. The encoder captures global information from the set of all particles, including the object’s shape and the historical motion of the particles, which are used to implicitly infer the object’s physical properties and dynamic state. This is essential for handling incomplete observations, where the feature encoder must extract occlusion-robust features for subsequent velocity decoding.
@@ III-B2 Neural Velocity Field
[…]
@@ III-B3 G2P
MPM [12] MSE $\downarrow$ $0.176_{\pm 0.107}$ $0.138_{\pm 0.072}$ $0.163_{\pm 0.148}$- $0.226_{\pm 0.026}$ $0.034_{\pm 0.014}$
GBND [58] $0.077_{\pm 0.033}$ $0.062_{\pm 0.025}$ $0.078_{\pm 0.028}$ $0.045_{\pm 0.008}$ $0.030_{\pm 0.011}$ $0.031_{\pm 0.014}$
Particle $0.059_{\pm 0.039}$ $0.061_{\pm 0.051}$ $0.060_{\pm 0.027}$ $0.025_{\pm 0.009}$ $0.021_{\pm 0.016}$ $0.038_{\pm 0.018}$
@@
MPM [12] CD $\downarrow$ $0.156_{\pm 0.091}$ $0.115_{\pm 0.054}$ $0.153_{\pm 0.207}$- $0.183_{\pm 0.032}$ $0.031_{\pm 0.010}$
GBND [58] $0.083_{\pm 0.034}$ $0.073_{\pm 0.027}$ $0.064_{\pm 0.016}$ $0.062_{\pm 0.014}$ $0.042_{\pm 0.007}$ $0.031_{\pm 0.013}$
Particle $0.051_{\pm 0.034}$ $0.059_{\pm 0.058}$ $0.043_{\pm 0.019}$ $0.032_{\pm 0.011}$ $0.025_{\pm 0.016}$ $0.044_{\pm 0.031}$
@@
MPM [12] EMD $\downarrow$ $0.093_{\pm 0.078}$ $0.081_{\pm 0.052}$ $0.092_{\pm 0.131}$- $0.110_{\pm 0.027}$ $0.021_{\pm 0.009}$
GBND [58] $0.035_{\pm 0.017}$ $0.036_{\pm 0.016}$ $0.032_{\pm 0.010}$ $0.032_{\pm 0.008}$ $0.016_{\pm 0.005}$ $0.016_{\pm 0.008}$
Particle $0.029_{\pm 0.024}$ $0.036_{\pm 0.037}$ $0.025_{\pm 0.013}$ $0.018_{\pm 0.006}$ $0.012_{\pm 0.010}$ $0.023_{\pm 0.016}$
@@
TABLE I: Quantitative Results on Dynamics Prediction. We compare our method with the Material Point Method (MPM) [12], Graph-Based Neural Dynamics (GBND) [58], and a particle-based dynamics model without the grid representation. We report the mean and standard deviation of the prediction error over a 3-second future horizon. The best results are highlighted in bold and blue.
@@ III-C Data Collection and Training
We collect training data through teleoperation and automatic annotation using foundation models. Specifically, we record multi-view RGB-D videos of random robot-object interactions. For each camera view, we apply Segment-Anything [18, 39] to extract persistent object masks across the video. The segmented objects are then cropped and tracked over time using CoTracker [15], providing 2D trajectories. Using depth information, we perform inverse projection to map these 2D velocities into 3D, resulting in multi-view fused point clouds with persistent particle tracking.
@@ III-D Rendering and Action-Conditioned Video Prediction
where $\mathbf{X}_{\text{GS}}$, $\mathbf{C}$, $\mathbf{R}_{\text{GS}}$, $\mathbf{S}$, and $\mathbf{O}$ represent the Gaussian kernels’ center location, color, rotation, scale, and opacity, respectively. To transform Gaussians between frames, we first apply the dynamics model to the point cloud set $\mathbf{X}$, yielding the next-frame prediction $\mathbf{\hat{X}}$. The points $\mathbf{\hat{X}}$ can either be sampled from $\mathbf{X}_{\text{GS}}$ or obtained from additional point cloud observations within the same coordinate frame with $\mathbf{X}_{\text{GS}}$. The 6-DoF motions of the Gaussian kernels are interpolated using Linear Blend Skinning (LBS) [45], which updates $\mathbf{X}_{\text{GS}}$ and $\mathbf{R}_{\text{GS}}$ by treating $\mathbf{X}$ as control points and interpolating their predicted motion to generate new Gaussian centers and rotations. We assume that the color, scale, and opacity of the Gaussian splatting remain constant.

### v2

@@ Abstract
Modeling the dynamics of deformable objects is challenging due to their diverse physical properties and the difficulty of estimating states from limited visual information. We address these challenges with a neural dynamics framework that combines object particles and spatial grids in a hybrid representation. Our particle-grid model captures global shape and motion information while predicting dense particle movements, enabling the modeling of objects with varied shapes and materials. Particles represent object shapes, while the spatial grid discretizes the 3D space to ensure spatial continuity and enhance learning efficiency. Coupled with Gaussian Splattings for visual rendering, our framework achieves a fully learning-based digital twin of deformable objects and generates 3D action-conditioned videos. Through experiments, we demonstrate that our model learns the dynamics of diverse objects–such as ropes, cloths, stuffed animals, and paper bags–from sparse-view RGB-D recordings of robot-object interactions, while also generalizing at the category level to unseen instances. Our approach outperforms state-of-the-art learning-based and physics-based simulators, particularly in scenarios with limited camera views. Furthermore, we showcase the utility of our learned models in model-based planning, enabling goal-conditioned object manipulation across a range of tasks. The project page is available at https://kywind.github.io/pgnd.
@@ I Introduction
Learning predictive models is crucial for a wide range of robotic tasks. In deformable object manipulation, an accurate predictive object dynamics model enables model-based planning, policy evaluation, and real-to-sim asset generation. However, developing dynamics models for deformable objects that are both accurate and generalizable remains a significant challenge. For example, physics-based simulators [hu2018moving, pbd2014] often struggle to generalize to the real world due to the inherent sim-to-real gap and the difficulties of system identification and state estimation. Meanwhile, video-based predictive models [finn2017deep, yang2023learning] are computationally expensive, lack 3D spatial understanding, and are highly sensitive to viewpoint and appearance changes.
@@
Recent advancements in deformable object modeling have focused on learning dynamics models for particles that encode 3D geometric information directly obtained from RGB-D cameras. The current state-of-the-art methods in this area utilize Graph Neural Networks (GNNs), where particle sets are modeled as spatial adjacency graphs, and message passing is performed to aggregate information and generate motion predictions [whitney2023learning3dparticlebasedsimulators, zhang2024dynamic]. However, these methods face significant challenges: the effectiveness of message passing is highly sensitive to the spatial distribution and connectivity of the graph nodes, making them vulnerable to partial observations. Additionally, insufficient message-passing steps may fail to capture global information, while excessive steps can lead to overly smoothed predictions. As a result, these approaches are often limited to relatively simple simulated environments or real-world objects with easily perceivable geometries, where graph construction is straightforward with nearest neighbors.
@@
To address these limitations, we introduce a novel class of dynamic models called particle-grid neural dynamics. This model leverages a hybrid representation that combines object particles with fixed spatial grids. It takes the kinematic states of the particles as input and predicts a spatial velocity field at fixed grid points. A global point cloud encoder is employed to capture comprehensive information from the particle set, effectively overcoming the connectivity challenges commonly faced by GNNs and enhancing the model’s robustness to incomplete observations. Additionally, this encoder allows for the processing of denser particle sets compared to sparse graph nodes, thereby enriching the model with finer geometric details. The grid representation further acts as a regularizer, ensuring spatial continuity in velocity predictions while also reducing computational costs when handling large numbers of particles. By combining object particles with spatial grids, our framework parameterizes dynamics in both Lagrangian and Eulerian coordinates, drawing an analogy to physics-based deformable object simulation methods [sulsky1995application, hu2018moving]. By utilizing neural networks as message integrators to bridge particle and grid representations, our model handles a diverse range of materials without requiring material-specific physical laws, and can operate effectively with only partial observations as input.
@@
Notably, our model is trained purely on RGB-D videos of robots interacting with objects through diverse and non-task-specific behaviors. To extract object particles from raw recordings, we have developed a 3D fusion and tracking framework. This framework utilizes foundational vision models [kirillov2023segment, ren2024grounded, karaev24cotracker3] to estimate segmentation masks and pixel tracks, which are then fused in 3D to generate persistent, dense 3D tracks that serve as training data for the neural dynamics model.
@@
In our experiments, we demonstrate that the proposed particle-grid neural dynamics model can effectively simulate a diverse range of challenging deformable objects, including ropes, cloth, plush toys, boxes, paper bags, and bread. We compare our model with existing state-of-the-art approaches and show that it consistently outperforms the baselines in prediction accuracy. Additionally, we highlight that our dynamics model seamlessly integrates with 3D appearance reconstruction methods such as 3D Gaussian Splatting (3DGS) [kerbl3Dgaussians], enabling the fusion of both techniques to generate highly realistic renderings of predictions. This improved accuracy in dynamics prediction directly translates to higher-fidelity renderings compared to existing methods. To further validate its robustness, we conduct experiments that demonstrate the model’s effectiveness under sparse-view conditions. Finally, we showcase the model’s applicability to deformable object manipulation by incorporating it into a Model Predictive Control (MPC) framework, underscoring its potential for real-world robotic applications.
@@ II Related Work
The simulation of deformable objects is essential for advancing both the modeling and robotic manipulation of soft, flexible materials. Researchers have introduced various analytical physics-based approaches, including, but not limited to, the mass-spring system [fastsimmassspring2013, diffclothsim2019], the Finite Element Method (FEM) [FEMdeform2006, peng20233dforcecontactestimation, du2021_diffpd], the Discrete Elastic Rods (DER) [DER2008, chen2024differentiableDERdlo], the Position-Based Dynamics (PBD) [pbd2014, corl2020softgym, pdbrope2023liu], and the Material Point Method (MPM) [sulsky1995application, hu2018moving, difftaichi2020, ma2023learning]. However, real-world analytical deformable modeling remains challenging due to the difficulty in property identification and state estimation. While our method uses a hybrid particle-grid representation similar to MPM, we leverage neural networks as message integrators and reduce dependence on full-state information as input. This enhances robustness to partial observations and enables the handling of a wide variety of deformable objects without requiring predefined material-specific constitutive laws. Recently, the emergence of 3D Gaussian Splatting (3DGS)  [kerbl3Dgaussians, luiten2023dynamic, duisterhof2023md] has enabled the fusion of Gaussian Splatting reconstructions with physics-based models to simulate the dynamics of deformable objects [xie2023physgaussian, qiu-2024-featuresplatting, springmassgs2024zhong, zhang2024physdreamer, abou-chakra2024physicallyembodiedgs]. In our particle-grid neural dynamics model, the particle motion predictions can also be integrated with Gaussian Splatting reconstructions, facilitating the simultaneous modeling and rendering of deformable objects, purely learned from real data.
@@
Learning-based dynamics models, which use deep neural networks to model the future evolution of dynamical systems, have demonstrated effectiveness across various robotic tasks [hoque2020visuospatial, learningdlo2020, kipf2019contrastive, zhou2024dinowmworldmodelspretrained, xue2023neural, bauer2024doughnet]. Among these learning-based methods, Graph-Based Neural Dynamics (GBND) have shown great promise, as they explicitly model spatial relational biases within complex physical systems [li2018learning, pfaff2020learning, sanchez2020learning, gnnrigidcontactdynamics2023allen, whitney2023learning3dparticlebasedsimulators]. Previous research has investigated the use of GBND across a range of material types, such as rigid bodies [huang2023defgraspnets, liu2023modelbased, ai2024robopack], plasticine [shi2022robocraft, shi2023robocook], fabrics [lin2022learning, longhini2023edo, puthuveetil2023robust, longhini2024clothsplatting], ropes [gnncable2022], and granular piles [Wang-RSS-23, shen2024babndlonghorizonmotionplanning]. Beyond simulation and single-material scenarios, GBND has also demonstrated flexibility and generalization in modeling diverse materials using a unified framework [zhang2024adaptigraph, zhang2024dynamic]. However, these approaches often operate on spatially sparse graph vertices, rely on expert knowledge to determine graph connectivity, and do not consider partial observations. For learning dense particle dynamics, whitney2024modelingrealworldhighdensity propose transformer-based backbones for higher computational efficiency. However, their work mainly focuses on modeling rigid objects for grasping and pushing tasks. In contrast, our work emphasizes deformable object modeling using a hybrid particle-grid neural dynamics framework, achieving dense particle prediction while remaining robust to incomplete observations and flexible for diverse types of deformable objects with distinct physical properties.
@@ III-A4 The Particle-Grid Model
Learning the neural dynamics parameters $\theta$ with an end-to-end neural network usually leads to unstable predictions due to accumulative error. Motivated by the use of hybrid Lagrangian-Eulerian representations in MPM [sulsky1995application], our model also utilizes a hybrid particle-grid representation to inject inductive bias related to spatial continuity and local information integration. Specifically, we define a uniformly distributed grid in the space as
@@
The parameters $l_{x},l_{y},l_{z},\delta$ control the spatial limits and resolution of the grid. Empirically, we set $l_{x},l_{y},l_{z}$ to 100 or 50 and $\delta$ to $1\text{\,}\mathrm{cm}$ or $2\text{\,}\mathrm{cm}$ to balance computational cost and resolution. To enforce translational invariance during prediction, we always translate the positions of the particles and the robot end effector to the volume defined by $\mathbf{G}_{l_{x},l_{y},l_{z},\delta}$.
@@ III-B1 Point Encoder
We use PointNet [qi2017pointnet] as the encoder for its efficiency and strong performance in extracting 3D point features. The encoder captures global information from the set of all particles, including the object’s shape and the historical motion of the particles, which are used to implicitly infer the object’s physical properties and dynamic state. This is essential for handling incomplete observations, where the feature encoder must extract occlusion-robust features for subsequent velocity decoding.
@@ III-B2 Neural Velocity Field
where $\mathcal{N}_{r}(\mathbf{X}_{t},\mathbf{x}_{g})$ is the set of indices of particles within $\mathbf{X}_{t}$ whose positions are within radius $r$ of the grid location $\mathbf{x}_{g}$. By incorporating the radius hyperparameter $r$, we can control the number of particle features a grid point attends to, thus encouraging the network to predict velocities that are dependent on local geometry. Empirically, we set $r=$0.2\text{\,}\mathrm{m}$\$.
@@ III-B3 G2P
MPM [hu2018moving] MDE $\downarrow$ $0.176_{\pm 0.107}$ $0.138_{\pm 0.072}$ $0.163_{\pm 0.148}$ - $226_{\pm 0.026} 0.0$$034_{\pm 0.014}$
GBND [zhang2024dynamic] $077_{\pm 0.033} 0.0$$62_{\pm 0.025} 0.0$$78_{\pm 0.028} 0.0$$45_{\pm 0.008} 0.0$$30_{\pm 0.011} 0.0$$031_{\pm 0.014}$
Particle $059_{\pm 0.039} 0.0$$61_{\pm 0.051} 0.0$$60_{\pm 0.027} 0.0$$25_{\pm 0.009} 0.0$$21_{\pm 0.016} 0.0$$038_{\pm 0.018}$
@@
MPM [hu2018moving] CD $\downarrow$ $0.156_{\pm 0.091}$ $0.115_{\pm 0.054}$ $0.153_{\pm 0.207}$ - $183_{\pm 0.032} 0.0$$031_{\pm 0.010}$
GBND [zhang2024dynamic] $083_{\pm 0.034} 0.0$$73_{\pm 0.027} 0.0$$64_{\pm 0.016} 0.0$$62_{\pm 0.014} 0.0$$42_{\pm 0.007} 0.0$$031_{\pm 0.013}$
Particle $051_{\pm 0.034} 0.0$$59_{\pm 0.058} 0.0$$43_{\pm 0.019} 0.0$$32_{\pm 0.011} 0.0$$25_{\pm 0.016} 0.0$$044_{\pm 0.031}$
@@
MPM [hu2018moving] EMD $\downarrow$ $093_{\pm 0.078} 0.0$$81_{\pm 0.052} 0.0$$092_{\pm 0.131} -$ $110_{\pm 0.027} 0.0$$021_{\pm 0.009}$
GBND [zhang2024dynamic] $035_{\pm 0.017} 0.0$$36_{\pm 0.016} 0.0$$32_{\pm 0.010} 0.0$$32_{\pm 0.008} 0.0$$16_{\pm 0.005} 0.0$$016_{\pm 0.008}$
Particle $029_{\pm 0.024} 0.0$$36_{\pm 0.037} 0.0$$25_{\pm 0.013} 0.0$$18_{\pm 0.006} 0.0$$12_{\pm 0.010} 0.0$$023_{\pm 0.016}$
@@
TABLE I: Quantitative Results on Dynamics Prediction. We compare our method with the Material Point Method (MPM) [hu2018moving], Graph-Based Neural Dynamics (GBND) [zhang2024dynamic], and a particle-based dynamics model without the grid representation. We report the mean and standard deviation of the prediction error over a 3-second future horizon. The best results are highlighted in bold and blue.
@@ III-C Data Collection and Training
We collect training data through teleoperation and automatic annotation using foundation models. Specifically, we record multi-view RGB-D videos of random robot-object interactions. For each camera view, we apply Segment-Anything [kirillov2023segment, ren2024grounded] to extract persistent object masks across the video. The segmented objects are then cropped and tracked over time using CoTracker [karaev24cotracker3], providing 2D trajectories. Using depth information, we perform inverse projection to map these 2D velocities into 3D, resulting in multi-view fused point clouds with persistent particle tracking.
@@ III-D Rendering and Action-Conditioned Video Prediction
where $\mathbf{X}_{\text{GS}}$, $\mathbf{C}$, $\mathbf{R}_{\text{GS}}$, $\mathbf{S}$, and $\mathbf{O}$ represent the Gaussian kernels’ center location, color, rotation, scale, and opacity, respectively. To transform Gaussians between frames, we first apply the dynamics model to the point cloud set $\mathbf{X}$, yielding the next-frame prediction $\mathbf{\hat{X}}$. The points $\mathbf{\hat{X}}$ can either be sampled from $\mathbf{X}_{\text{GS}}$ or obtained from additional point cloud observations within the same coordinate frame with $\mathbf{X}_{\text{GS}}$. The 6-DoF motions of the Gaussian kernels are interpolated using Linear Blend Skinning (LBS) [10.1145/1276377.1276478], which updates $\mathbf{X}_{\text{GS}}$ and $\mathbf{R}_{\text{GS}}$ by treating $\mathbf{X}$ as control points and interpolating their predicted motion to generate new Gaussian centers and rotations. We assume that the color, scale, and opacity of the Gaussian splatting remain constant.